# Symmetry-enriched topological order and quasi-fractonic behavior in $\mathbb{Z}_{N}$ stabilizer codes

### v1

@@ VI.2 DCY model
We believe this is the first accurate calculation of the GSD of the DCY model.
VI.3 Product code
We obtain the anyon types and their fusion rules for the $\mathbb{Z}_{N}$ qudit product code specified by univariate polynomials $f(x)$ and $g(y)$ in the following. According to Theorem 1, the product code is topological if and only if $fg=0$ and $f+g$ is not a monomial over $\mathbb{Z}_{p}[x^{\pm},y^{\pm}]$ for all primes $p\mid N$. For prime qudit dimensions, the fusion rules of product codes can be easily obtained by long division. Consequently, the fusion rules of product codes with an arbitrary qudit dimension $N$ can be directly obtained using Theorem 2. The result is given by
@@ VI.3 Product code
VI.4 BB code with toric layouts
Figure 5: The topological index of the $\mathbf{BB}_{p}(f,g)$ code with $f$ and $g$ from Eq. (55) for two special cases: (a) all coefficients are non-zero modulo $p$. (b) only $d^{(1)}$ modulo $p$ is zero. For both cases, the parameter space is divided into five regimes by four hyperbolas, and the corresponding expression of the topological index in each regime is indicated. Parameters are restricted to $\beta\geq\alpha\geq 0$ for (a), and to $\alpha,\beta\geq 0$ and $a\leq 0$ for (b), as all other values can be reduced to this range by suitable origin shifts and axis reflections.
@@ VI.4 BB code with toric layouts
$\mathcal{C}_{p^{k}}\cong\mathbb{Z}_{p^{k}}^{2Q}$ (57)
@@ VII Computational algebraic methods
4 $1+x+y^{2}$ $1+y+x^{3}$ $\mathbb{Z}_{4}^{8}$ $l_{x}=10,l_{y}=30$ $\langle x^{10},x^{6}y^{6}\rangle,\langle y^{30},x^{12}y^{2}\rangle$ 11
@@
4 $x^{5}y^{-1}+y+1$ $y^{5}x^{-1}+x+1$ $\mathbb{Z}_{4}^{24}$ $l_{x}=l_{y}=19530$ $\langle x^{19530},x^{12090}y^{930}\rangle,\langle y^{19530},x^{930}y^{12090}\rangle$ 11
4 $2+x+xy^{2}$ $1+2y+x^{4}y^{3}$ $\mathbb{Z}_{4}^{16}$ $l_{x}=16,l_{y}=4$ $\langle x^{16},x^{8}y^{2}\rangle,\langle y^{4},x^{8}y^{2}\rangle$ 88
@@
8 $x^{3}+y^{2}+y$ $y^{3}+x^{2}+x$ $\mathbb{Z}_{8}^{16}$ $l_{x}=l_{y}=48$ $\langle x^{48},x^{24}y^{24}\rangle,\langle y^{48},x^{24}y^{24}\rangle$ $8^{5}$
9 $4+6x+x^{-1}y^{2}$ $7+5y+2x^{-3}y^{-4}$ $\mathbb{Z}_{9}^{6}$ $l_{x}=312,l_{y}=624$ $\langle x^{312},x^{306}y^{12}\rangle,\langle y^{624},x^{12}y^{600}\rangle$ $9^{6}$
@@
Table 1: Topological properties of various bivariate-bicycle (BB) codes, including their defining polynomials $f$ and $g$, fusion rules, periodicities ($l_{x},l_{y}$), mobility groups, and ground state degeneracies (GSD) computed for system size $L_{x}=L_{y}=18$. Here, $\langle\bullet,\bullet\rangle$ denotes the subgroup of the translation group $\Lambda$ generated by a pair of elements in $\Lambda$.
@@ VII.1.2 Crucial properties of Gröbner bases
$\{r|r=\sum_{\mathbf{w}}c_{\mathbf{w}}X^{\mathbf{w}},c_{\mathbf{w}}\in\mathbb{Z}_{N_{\mathbf{w}}}\},$ (63)
where $X^{\mathbf{w}}=x^{\omega_{x}}y^{\omega_{y}}t^{\omega_{t}}u^{\omega_{u}}$ denotes a monomial in the polynomial ring $\mathbb{Z}[x,y,t,u]$ with $\mathbf{w}=(\omega_{x},\omega_{y},\omega_{t},\omega_{u})\in\mathbb{Z}_{\geq 0}^{4}$. The range of the coefficient $c_{\mathbf{w}}$ is $\mathbb{Z}_{N_{\mathbf{w}}}$, with $N_{\mathbf{w}}$ defined as
$N_{\mathbf{w}}=\gcd(C_{\mathbf{w}}),\quad C_{\mathbf{w}}:=\{\operatorname{Lc}(h)|X^{\mathbf{w}}=0\bmod\operatorname{Lm}(h),h\in G\},$ (64)
and $\gcd(C_{\mathbf{w}})$ denotes the greatest common divisor of all elements in $C_{\mathbf{w}}$  [^2] If $C_{\mathbf{w}}=\varnothing$, then we take $N_{\mathbf{w}}=N$.. This is because the leading term $cX^{\mathbf{w}}$ of a polynomial can be reduced by $G$ through the division algorithm if and only if its coefficient lies in the ideal $(C_{\mathbf{w}})_{\mathbb{Z}}$.
@@ VII.2 Topological properties of BB codes from Gröbner basis
Moreover, the anyon periodicity and the mobility group of the $\mathbf{BB}_{N}(f,g)$ code are given by Eq. (27) and Eq. (28), respectively. Determining them requires testing whether $x^{\alpha}y^{\beta}-1$ with $\alpha,\beta\in\mathbb{Z}$ lies in $(f,g)$ in $R$. This is an ideal membership problem. Using the isomorphism in Eq. (61), it is equivalent to the ideal membership problem of I=(N,f,g,xt1,yu1)I=(N,f,g,xt-1,yu-1) in $\mathbb{Z}[x,y,t,u]$. Thus, the anyon periodicity and the mobility group are directly obtained by testing if $x^{\alpha}y^{\beta}-1$ reduces to zero in $I$ by the division algorithm with respect to a Gröbner basis $G$.
@@ VII.3 Examples
The ideal membership tests in Macaulay 2 show that the periodicities along the $x$-and $y$-directions are $l_{x}=2$ and $l_{y}=8$, and that the mobility group is $\langle x^{2},x^{4}y^{4}\rangle$, or equivalently $\langle y^{8},x^{4}y^{4}\rangle$. Finally, applying the same Gröbner basis methods to the ideal for a finite torus, we find that for the code of size $L_{x}=L_{y}=18$, the ground state degeneracy is 16.
The simple sample code of this example is provided in Appendix F, which also demonstrates how the topological condition can also be verified directly by definition. The topological properties of several other BB codes, determined using this method, are summarized in Table 1.
@@ VIII Conclusion
In this work, we explored several generalizations of the toric code model for qudits. By utilizing the polynomial representation, we systematically characterized essential topological properties, including the topological condition, the anyon types and their fusion rules of qudit BB codes. We established a convenient finiteness criterion for the topological condition of qudit BB codes. Our analysis revealed that both the topological condition and the fusion rules for general qudit dimensions can be understood by examining prime-dimensional BB codes. This insight allowed us to employ an algebraic-geometric framework, specifically leveraging the BKK theorem, to systematically evaluate the anyon types and their fusion rules. Additionally, we examined that unconventional phenomena in qudit BB codes, including the quasi-fractonic behavior and size-dependent ground state degeneracy, originate from their symmetry-enriched topological order. Furthermore, we analytically derived the key topological properties of the WCF and DCY models. Our work resolves the mobility puzzle and presents the first accurate analytical computation of the ground state degeneracy for the DCY model. Lastly, for more complicated BB codes, we proposed a computational algebraic method to calculate their topological properties based on the Gröbner bases over the ring of integers.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 12474145, No. 12522502, and No. 12447101).
@@ A.3 Proof of Corollary 1
$0\to R_{p}\xrightarrow{\partial_{Bp}}R_{p}^{2}\xrightarrow{\varepsilon_{Ap}}R_{p}$ (75)
is exact. (Injectivity of $\partial_{Bp}$ is trivial, as $R_{p}$ is an integral domain). We invoke the Buchsbaum-Eisenbud exactness criterion [97] for a complex of free modules over the ring $R_{p}$. The criterion requires, among other things, that the codimension (or height) of the ideal of maximal minors of $\varepsilon_{Ap}$ be at least 2. The matrix for $\varepsilon_{Ap}$ is $(f_{p},g_{p})$, so its ideal of maximal ($1\times 1$) minors is simply the ideal $I=(f_{p},g_{p})$. The ring $R_{p}$ has Krull dimension 2. Therefore, the condition becomes $\operatorname{codim}(f_{p},g_{p})\geq 2$. Since codimension cannot exceed the dimension of the ring, this forces $\operatorname{codim}(f_{p},g_{p})=2$.
@@ Appendix B Legitimate excitation patterns
In this section, we show that every element of $R^{2}$ is a legitimate excitation pattern, meaning that it can be created by either a local or an infinitely extended operator. By linearity, it is sufficient to prove this for the generators $(1,0)^{T}$ and $(0,1)^{T}$.
@@
Next, according to Fact 1 (given below), it is possible to choose a weighted degree such that the lowest-degree term in $r$ is unique; let’s call it $cx^{a_{1}}y^{a_{2}}$. The lowest-degree term of $r^{k}$ is then $c^{k}x^{ka_{1}}y^{ka_{2}}$. By construction, $p\nmid c$, so $c$ is invertible in $\mathbb{Z}_{p^{k}}$. Dividing both sides of Eq. (78) by $c^{k}x^{ka_{1}}y^{ka_{2}}$ gives
@@
2) Select a Unique Candidate. We use lexicographical order to find a unique “smallest” vector in $S$. A vector $\mathbf{a}$ is lexicographically smaller than $\mathbf{b}$ if the first non-zero entry in the difference vector $\mathbf{b}-\mathbf{a}$ is positive. Since $S$ is a finite, non-empty set of vectors, it has a unique lexicographically minimal element. Let’s call it $\mathbf{a}_{0}$.
@@ Appendix C Proof of Theorem 2
In this section, we provide a proof of Theorem 2.
@@
where $F_{i}$ denotes free $S$-modules, whereas $\varphi_{i}$ represents a module homomorphism from $F_{i+1}$ to $F_{i}$ for each $i$. Tensoring the free resolution Eq. (83) with $M$ induces the following chain complex:
@@
A proof can be found in Ref. [99, Tag 051H]. To compute the anyon types and their fusion rules, we consider $\mathcal{Q}=R/(f,g)$ for the $\mathbf{BB}_{p^{k}}(f,g)$ code. The following lemma shows that the $\operatorname{Tor}_{1}=0$ condition holds for topological $\mathbb{Z}_{p^{k}}$ BB codes.
@@
This corollary is proved by directly applying Lemma 4, where we take $S\coloneqq\mathbb{Z}_{p^{k}}$, $I\coloneqq(p)$, and $M\coloneqq\mathcal{Q}$. By definition, we have $M/IM=\mathcal{Q}_{p}$ and $S/I=\mathbb{Z}_{p}$. In this scenario, condition (i) is evident, condition (ii) holds since $\mathbb{Z}_{p}$ is a field, and condition (iii) is valid according to Lemma 86. Thus, Corollary 2 follows since $\mathcal{C}=\mathcal{Q}\oplus\overline{\mathcal{Q}}$. ∎
@@
$P_{A}\coloneqq\frac{1}{\left|S_{AL}\right|}\sum_{s_{A}\in S_{AL}}s_{A}\quad\text{and}\quad P_{B}\coloneqq\frac{1}{\left|S_{BL}\right|}\sum_{s_{B}\in S_{BL}}s_{B}.$ (94)

### v2

@@ VI.2.3 Ground state degeneracy
We believe that this is an accurate calculation of the GSD of the DCY model. It corrects the previous result reported in Ref. [54].
As an explicit check, consider the case $N=2$ and $a=1$ (so $f=x^{2}-1$ and $g=y^{2}-1$) on a torus with $L_{x}=L_{y}=l$. In this case, the model reduces to four decoupled toric codes supported on the four sublattices distinguished by the parities of $(i,j)$: (i) $i$ odd, $j$ odd; (ii) $i$ even, $j$ odd; (iii) $i$ odd, $j$ even; and (iv) $i$ even, $j$ even. When $l$ is even, periodic boundary conditions preserve these sublattices, yielding $\operatorname{GSD}=4^{4}=256$; when $l$ is odd, periodic boundary conditions identify them, leaving a single toric code with $\operatorname{GSD}=4$. Our Eq. (56) correctly captures this system-size dependence. In contrast, Ref. [54] [Eq. (27) therein] predicts $\operatorname{GSD}=256$ for this example independent of $l$, and thus misses the odd-$l$ behavior.
VI.3 Hypergraph product codes
We obtain the anyon types and their fusion rules for the $\mathbb{Z}_{N}$ qudit HGP code specified by univariate polynomials $f(x)$ and $g(y)$ in the following. According to Theorem 1, the HGP code is not topological if and only if $fg=0$ and $f+g$ is not a monomial over $\mathbb{Z}_{p}[x^{\pm},y^{\pm}]$ for all primes $p\mid N$. For prime qudit dimensions, the fusion rules of HGP codes can be easily obtained by long division. Consequently, the fusion rules of HGP codes with an arbitrary qudit dimension $N$ can be directly obtained using Theorem 2. The result is given by
@@ VI.3 Hypergraph product codes
VI.4 BB codes with toric layouts
Figure 5: The topological index of the $\mathbf{BB}_{p}(f,g)$ code with $f$ and $g$ from Eq. (58) for two special cases: (a) all coefficients are non-zero modulo $p$. (b) only $d^{(1)}$ modulo $p$ is zero. For both cases, the parameter space is divided into five regimes by four hyperbolas, and the corresponding expression of the topological index in each regime is indicated. Parameters are restricted to $\beta\geq\alpha\geq 0$ for (a), and to $\alpha,\beta\geq 0$ for (b), as all other values can be reduced to this range by suitable origin shifts and axis reflections.
@@ VI.4 BB codes with toric layouts
$\mathcal{C}_{p^{k}}\cong\mathbb{Z}_{p^{k}}^{2Q},$ (60)
@@ VII Computational algebraic methods
4 $1+x+y^{2}$ $1+y+x^{3}$ $\mathbb{Z}_{4}^{8}$ $l_{x}=10,l_{y}=30$ $\langle x^{10},x^{6}y^{6}\rangle,\langle y^{30},x^{2}y^{12}\rangle$ 11
@@
4 $x^{5}y^{-1}+y+1$ $y^{5}x^{-1}+x+1$ $\mathbb{Z}_{4}^{48}$ $l_{x}=l_{y}=19530$ $\langle x^{19530},x^{12090}y^{930}\rangle,\langle y^{19530},x^{930}y^{12090}\rangle$ 11
4 $2+x+xy^{2}$ $1+2y+x^{4}y^{3}$ $\mathbb{Z}_{4}^{16}$ $l_{x}=16,l_{y}=4$ $\langle x^{16},x^{8}y^{2}\rangle,\langle y^{4},x^{8}y^{2}\rangle$ $4^{3}$
@@
8 $x^{3}+y^{2}+y$ $y^{3}+x^{2}+x$ $\mathbb{Z}_{8}^{16}$ $l_{x}=l_{y}=48$ $\langle x^{48},x^{24}y^{24}\rangle,\langle y^{48},x^{24}y^{24}\rangle$ $8^{9}\times 2$
9 $4+6x+x^{-1}y^{2}$ $7+5y+2x^{-3}y^{-4}$ $\mathbb{Z}_{9}^{22}$ $l_{x}=39348,l_{y}=78696$ $\langle x^{39348},x^{39342}y^{12}\rangle,\langle y^{78696},x^{6}y^{78684}\rangle$ $9^{4}$
@@
Table 1: Topological properties of various $\mathbb{Z}_{N}$ bivariate-bicycle (BB) codes, including their defining polynomials $f$ and $g$, fusion rules, periodicities ($l_{x},l_{y}$), mobility groups, and ground state degeneracies (GSD) computed for system size $L_{x}=L_{y}=18$. Here, $\langle\bullet,\bullet\rangle$ denotes the subgroup of the translation group $\Lambda$ generated by a pair of elements in $\Lambda$.
@@ VII.1.2 Crucial properties of Gröbner bases
$\{r|r=\sum_{\mathbf{w}}c_{\mathbf{w}}\mathcal{X}^{\mathbf{w}},c_{\mathbf{w}}\in\mathbb{Z}_{N_{\mathbf{w}}}\},$ (66)
where $\mathcal{X}^{\mathbf{w}}=x^{\omega_{x}}y^{\omega_{y}}t^{\omega_{t}}u^{\omega_{u}}$ denotes a monomial in the polynomial ring $\mathbb{Z}[x,y,t,u]$ with $\mathbf{w}=(\omega_{x},\omega_{y},\omega_{t},\omega_{u})\in\mathbb{Z}_{\geq 0}^{4}$. The range of the coefficient $c_{\mathbf{w}}$ is $\mathbb{Z}_{N_{\mathbf{w}}}$, with $N_{\mathbf{w}}$ defined as
$N_{\mathbf{w}}=\gcd(C_{\mathbf{w}}),\quad C_{\mathbf{w}}:=\{\operatorname{Lc}(h)|\mathcal{X}^{\mathbf{w}}=0\bmod\operatorname{Lm}(h),h\in G\},$ (67)
and $\gcd(C_{\mathbf{w}})$ denotes the greatest common divisor of all elements in $C_{\mathbf{w}}$  [^2] If $C_{\mathbf{w}}=\varnothing$, then we take $N_{\mathbf{w}}=N$.. This is because the leading term $c\mathcal{X}^{\mathbf{w}}$ of a polynomial can be reduced by $G$ through the division algorithm if and only if its coefficient lies in the ideal $(C_{\mathbf{w}})_{\mathbb{Z}}$.
@@ VII.2 Topological properties of BB codes from Gröbner basis
Moreover, the anyon periodicity and the mobility group of the $\mathbf{BB}_{N}(f,g)$ code are given by Eqs. (27) and (28), respectively. Determining them requires testing whether $x^{\alpha}y^{\beta}-1$ with $\alpha,\beta\in\mathbb{Z}$ lies in $(f,g)$ in $R$. This is an ideal membership problem. Using the isomorphism in Eq. (64), it is equivalent to the ideal membership problem of I=(N,f,g,xt1,yu1)I=(N,f,g,xt-1,yu-1) in $\mathbb{Z}[x,y,t,u]$. Thus, the anyon periodicity and the mobility group are directly obtained by testing if $x^{\alpha}y^{\beta}-1$ reduces to zero in $I$ by the division algorithm with respect to a Gröbner basis $G$.
@@ VII.3 Examples
The ideal membership tests in Macaulay 2 show that the periodicities along the $\hat{x}$ and $\hat{y}$ directions are $l_{x}=2$ and $l_{y}=8$, and that the mobility group is $\langle x^{2},xy^{4}\rangle$, or equivalently $\langle y^{8},xy^{4}\rangle$. Finally, applying the same Gröbner basis methods to the ideal for a finite torus, we find that for the code of size $L_{x}=L_{y}=18$, the ground state degeneracy is 16.
The simple sample code of this example is provided in Appendix F, which also demonstrates how the topological condition can be verified directly by definition. The topological properties of several other BB codes determined using this method are summarized in Table 1.
@@ VIII Conclusion
In this work, we explored several generalizations of the toric code model for qudits. By utilizing the polynomial representation, we systematically characterized essential topological properties, including the topological condition, the anyon types and their fusion rules of qudit BB codes. We established a convenient finiteness criterion for the topological condition of qudit BB codes. Our analysis revealed that both the topological condition and the fusion rules for general qudit dimensions can be understood by examining prime-dimensional BB codes. This insight allowed us to employ an algebraic-geometric framework, specifically leveraging the BKK theorem, to systematically evaluate the anyon types and their fusion rules. Additionally, we examined that unconventional phenomena in qudit BB codes, including the quasifractonic behavior and size-dependent ground state degeneracy, originate from their symmetry-enriched topological order. Furthermore, we analytically derived the key topological properties of the WCF and DCY models. Our work resolves the mobility puzzle and presents an accurate analytical computation of the ground state degeneracy for the DCY model. Lastly, for more complicated BB codes, we proposed a computational algebraic method to calculate their topological properties based on the Gröbner bases over the ring of integers. We hope that our framework will facilitate future work on broader classes of qudit stabilizer codes and their topological and symmetry-enriched properties.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 12522502, No. 12474145, No. 12275331, and No. 12447101), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB1680000), the Fundamental and Interdisciplinary Frontier Research Priority Program of Chinese Academy of Sciences (Grant No. XDB0920000), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301602).
@@ A.3 Proof of Corollary 1
$R_{p}\xrightarrow{\partial_{Bp}}R_{p}^{2}\xrightarrow{\varepsilon_{Ap}}R_{p}$ (78)
is exact, so the chain complex $0\to R_{p}\xrightarrow{\partial_{Bp}}R_{p}^{2}\xrightarrow{\varepsilon_{Ap}}R_{p}$ (the injectivity of $\partial_{Bp}$ is trivial, as in Lemma 2). We invoke the Buchsbaum-Eisenbud exactness criterion [117] for a complex of free modules over the ring $R_{p}$. The criterion requires, among other things, that the codimension (or height) of the ideal of maximal minors of $\varepsilon_{Ap}$ be at least 2. The matrix for $\varepsilon_{Ap}$ is $(f_{p},g_{p})$, so its ideal of maximal ($1\times 1$) minors is simply the ideal $I=(f_{p},g_{p})$. The ring $R_{p}$ has Krull dimension 2. Therefore, the condition becomes $\operatorname{codim}(f_{p},g_{p})\geq 2$. Since codimension cannot exceed the dimension of the ring, this forces $\operatorname{codim}(f_{p},g_{p})=2$.
@@ Appendix B Legitimate excitation patterns
In this appendix, we show that every element of $R^{2}$ is a legitimate excitation pattern, meaning that it can be created by either a local or an infinitely extended operator. By linearity, it is sufficient to prove this for the generators $(1,0)^{T}$ and $(0,1)^{T}$.
@@
Next, according to Fact 1 (given below), it is possible to choose a weighted degree such that the lowest-degree term in $r$ is unique; let us call it $cx^{a_{1}}y^{a_{2}}$. The lowest-degree term of $r^{k}$ is then $c^{k}x^{ka_{1}}y^{ka_{2}}$. By construction, $p\nmid c$, so $c$ is invertible in $\mathbb{Z}_{p^{k}}$. Dividing both sides of Eq. (81) by $c^{k}x^{ka_{1}}y^{ka_{2}}$ gives
@@
2) Select a unique candidate. We use lexicographical order to find a unique “smallest” vector in $S$. A vector $\mathbf{a}$ is lexicographically smaller than $\mathbf{b}$ if the first non-zero entry in the difference vector $\mathbf{b}-\mathbf{a}$ is positive. Since $S$ is a finite, non-empty set of vectors, it has a unique lexicographically minimal element. Let us call it $\mathbf{a}_{0}$.
@@ Appendix C Proof of Theorem 2
In this appendix, we provide a proof of Theorem 2.
@@
where $F_{i}$ denotes free $S$-modules, whereas $\varphi_{i}$ represents a module homomorphism from $F_{i}$ to $F_{i-1}$ for each $i$. Tensoring the free resolution Eq. (86) with $M$ induces the following chain complex:
@@
A proof can be found in Ref. [119, Tag 051H]. To compute the anyon types and their fusion rules, we consider $\mathcal{Q}=R/(f,g)$ for the $\mathbf{BB}_{p^{k}}(f,g)$ code. The following lemma shows that the $\operatorname{Tor}_{1}=0$ condition holds for topological $\mathbb{Z}_{p^{k}}$ BB codes.
@@
This corollary is proved by directly applying Lemma 4, where we take $S\coloneqq\mathbb{Z}_{p^{k}}$, $I\coloneqq(p)$, and $M\coloneqq\mathcal{Q}$. By definition, we have $M/IM=\mathcal{Q}_{p}$ and $S/I=\mathbb{Z}_{p}$. In this scenario, condition (i) is evident, condition (ii) holds since $\mathbb{Z}_{p}$ is a field, and condition (iii) is valid according to Lemma 89. Thus, Proposition 1 follows since $\mathcal{C}=\mathcal{Q}\oplus\overline{\mathcal{Q}}$. ∎
@@
$P_{A}\coloneqq\frac{1}{\left|S_{A}\right|}\sum_{s_{A}\in S_{A}}s_{A}\quad\text{and}\quad P_{B}\coloneqq\frac{1}{\left|S_{B}\right|}\sum_{s_{B}\in S_{B}}s_{B}.$ (97)